PLEASE help me on this question. It is on a take home test that is due TOMMOROW.
2. Give two solutions to the equation: determinant:
1 x y
-1 2 -2 =0
2 4 6
Thank you very much

Answer by stanbon(75887)

Give two solutions to the equation: determinant:
1 x y
-1 2 -2 =0
2 4 6
[12-4y-4x]-[4y-8-6x]=0
20-8y+2x=0
10-4y+x=0
4y=x+10
y=(1/4)x+(5/2)
Let x=0 then y=5/2
Let x=4 then y= 7/2
Cheers,
Stan H.